OpenStudy (anonymous):

Max-Min Problem help please 18. A cylinder storage tank is to contain V=16,000pi cubic feet (about 400,00 gallons). the cost of the tank is proportional to its area, so the minimal-cost tank will be the one with minimum area. The volume (V) of a cylinder of radius r and height h is (pi)(r^2)(h). its surface area is the area of top and bottom, 2(pi)(r^2), plus the side area 2(pi)(r^2)(h). find the dimensions, r and h, of the minimal-area tank.

OpenStudy (ranga):

The side area of the cylinder should be 2(pi)rh and not 2(pi)(r^2)(h)

OpenStudy (anonymous):

yes sorry about that

OpenStudy (ranga):

V = (pi)(r^2)(h) = 16,000 ft^3 h = 16000 / ((pi)r^2) --- (1) A = 2(pi)r^2 + 2(pi)rh = 2(pi)(r^2 + rh) ---- (2) Put h = 6000 / ((pi)r^2) from (1) into (2). You will have A as a function of just r (no h). Find dA/dr, equate to 0 and solve for r. Then put r in (1) and solve for h.

OpenStudy (ranga):

It should be: Put h = 16000 / ((pi)r^2) from (1) into (2).

OpenStudy (ranga):

Can you take it from here?

OpenStudy (ranga):

Yes, A = 2(pi)r^2+32000/r Are you in calculus or pre-calculus?

OpenStudy (ranga):

DO you know how to find the derivative? If not, are you allowed to use graphing calculator?

OpenStudy (ranga):

I am not sure what they expect you to use to minimize area A in the function: A = 2(pi)r^2+32000/r

OpenStudy (ranga):

With calculus or graphing calculator, it is straightforward.

OpenStudy (anonymous):

well graphing calculator will only be allowed 15 mins of the test

OpenStudy (ranga):

If you use graphing calculator, then simply plot the function: y = 2(pi)x^2+32000/x and find the lowest point on the curve. (we are replacing r with x so it will be acceptable to some calculators). The x value of the lowest point on the curve will be the radius that will minimize the area A. I plotted it and I got r = 13.66 feet. Put this r value in (1) and find h.

OpenStudy (anonymous):

27.29

OpenStudy (ranga):

Yes, radius = 13.66 ft and height = 27.29 ft. will have the minimum surface area. When plotting the curve we need to choose the proper scale factor. I chose to plot x between 0 and 40 and y between 0 and 32000. And I got a good graph where the graphing calculator put a dot on the minimum point on the curve from which I was able to find x (or radius) = 13.66 ft.

OpenStudy (ranga):

I am answering one more question so I will show up when I am done there.

OpenStudy (anonymous):

ok no problem

OpenStudy (ranga):

V = (pi)(r^2)(h) = 16,000 ft^3 h = 16000 / ((pi)r^2) --- (1) Now there is no top. So the area A = (pi)r^2 + 2(pi)rh Cost C = 10(pi)r^2 + 8.64 * 2(pi)rh C = 10(pi)r^2 + 17.28(pi)rh ---- (2) Put h from (1) into (2) We need to minimize C. Plot C as a function r. (or make C as y and r as x). Find for which value of x (or r) the cost curve attains a minimum. Then put the value of r in (1) and find h.

OpenStudy (anonymous):

10(pi)(r^2)+276480/r

OpenStudy (ranga):

C = 31.415r^2 + 276480/r Can you plot this on your graphing calculator and tell me what value of x gives the lowest point on the curve?

OpenStudy (ranga):

I get radius r = 16.39 feet when C will be the lowest. The big number you got is the y-coordinate or the C value and they are not asking for that. So put r = 16.39 in (1) and find h.

OpenStudy (ranga):

Yes. r = 16.39 feet and h = 18.96 feet will minimize the cost.

OpenStudy (anonymous):

thank u so much

OpenStudy (ranga):

you are welcome.